OpenStudy (anonymous):

Let h(x)=xg(x), where g(x)=the inverse of f(x). Find h'(5) using the chart.....(type that in a minute)

OpenStudy (anonymous):

x: 2, 3, 5 f(x): 4, 5, 1 f'(x): -1, 2, 3

OpenStudy (mathmate):

Let's make the table for g(x)=f-1(x) x = 4 5 1 g(x)=2 3 5 h(x)=xg(x) h'(x)=g(x)+xg'(x) [product rule] Apply f'^1(x)=1/f'(f^-1(x)) g'(x)=1/(f'(g(x))) g'(5)=1/f'(g(5)) =1/f'(3) =1/2 Substitute in above h'(5)=g(5)+5g'(5) =3+5*(1/2) =11/2